OpenStudy (anonymous):

Can someone graph this for me -4x + 3y = 6.

OpenStudy (anonymous):

you can make this easier by converting it to the equation, y=2+(4/3)x

OpenStudy (anonymous):

okay so like find the slope and all?

OpenStudy (anonymous):

yeah, so based on the form y=mx+b, m (the slope)= 4/3

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you so much!!